OpenStudy (anonymous):

Solve the equation by completing the square. x2 − 2x − 2 = 0

OpenStudy (anonymous):

take half the coefficent of the b term or the middle term. In this case it is -2 so take half of that. You should get -1. Now foil this out: (x-1)^2

OpenStudy (anonymous):

you should get: x^2-2x+1

OpenStudy (anonymous):

caught my own error there

OpenStudy (anonymous):

you will get: (x-1)^2-1-2=0, then evlluate this and get: (x-1)^2=3

OpenStudy (dumbcow):

yes :)

OpenStudy (anonymous):

take the square root of both sides: x-1=sqrt3

OpenStudy (anonymous):

should i say plus or minus sqrt 3

OpenStudy (anonymous):

now, you add 1 to both sides and get your answer as being: x=1 plus or minus sqrt3

OpenStudy (anonymous):

Thank you so much!!